belloki:

Harry is trying to solve the equation y = 2x2 − x − 6 using the quadratic formula. He has made an error in one of the steps below. Find the step where Harry went wrong. (1 point) Step 1: x equals the negative of negative 1 plus or minus the square root of the quantity negative one squared minus 4 times 2 times negative six, end quantity, all over 2 times 2. Step 2: x equals the negative of negative 1 plus or minus the square root of negative one plus forty-eight all over two times 2. Step 3: x equals the negative of negative 1 plus or minus the square root of forty-seven all over two times 2. Step 4: x equals 1 plus or minus the square root of forty-seven all over 4. Step 1 Step 2 Step 3 Step 4

Mercury:

thoughts? as a reminder, the quadratic formula is: \[\frac{ -b±\sqrt{b^{2}-4ac} }{ 2a }\] where a, b, and c correspond to ax^2 + bx + c in your case, since the equation is 2x^2 - x - 6, a = 2, b = -1, c = -6 substitute these into the equation, and follow along with his steps. as a hint, he made a sign error somewhere